OpenStudy (anonymous):

if f(x)=x3 and g(x)=x-1 which of the following statements is true A (g*f)(0) =3 B (f*g)(0)=1 C (f*g)(2)=1 D (g*f)(2)=6

OpenStudy (mathmate):

(g*f)(0)=g(f(0))=g(0^3)=g(0)=0-1=-1 \( \ne \) 3 You can do similarly for other cases to get the correct answer.

OpenStudy (anonymous):

its x qube i couldnt get it to do that when i put in the question

OpenStudy (mathmate):

We usually put x^3 to mean x cubed, and so on, but I understood what it meant. :)

OpenStudy (anonymous):

okey thanks